OpenStudy (anonymous):

Solve 2x2 + 3x + 5 = 0. Round solutions to the nearest hundredth I got -4.75 and -1.25

OpenStudy (anonymous):

need to make sure I solved correctly

OpenStudy (anonymous):

\[ 2x^2 + 3x + 5 = 0\\ b^2 -4 a c=3^2 -4(2)(5)=9 -40=-31\\ \] So the two roots are complex.

OpenStudy (anonymous):

\[ \begin{array}{c} x=\frac{1}{4} \left(-3-i \sqrt{31}\right) \\ x=\frac{1}{4} \left(-3+i \sqrt{31}\right) \\ \end{array} \]

OpenStudy (anonymous):

So there is no real solution?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Did you type your problem correctly?

OpenStudy (anonymous):

2x^2+3x+5=0

OpenStudy (anonymous):

May be they want you to write the roots as \[ \begin{array}{c} x=-0.75-1.39194 i \\ x=-0.75+1.39194 i \\ \end{array} \]

OpenStudy (anonymous):

To the nearest hundredth \[ \begin{array}{c} x=-0.75-1.39 i \\ x=-0.75+1.39 i \\ \end{array} \]

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

Yes I'm just confused as to how you got that answer

OpenStudy (anonymous):

Use the quadratic formula and your Calculator.

OpenStudy (anonymous):

Oh I did +4ac instead of -4ac

OpenStudy (anonymous):

A big difference.

OpenStudy (anonymous):

I'm getting a -31 and you can't square - numbers

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[ \sqrt{-31} = i \sqrt{31} \]

OpenStudy (anonymous):

Where \[ i^2 = -1 \]

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

Do you know how to use the quadratic formula? Do you know what a complex number is?

OpenStudy (anonymous):

quadratic formula , yes complex number, no

OpenStudy (anonymous):

For quadratic formula, you need to know complex numbers in case b^2 - 4 ac <0

OpenStudy (anonymous):

Okay and what if you do have a complex number?